# Toward Generalist Semi-supervised Regression via Decoupled Representation Distillation

### v1

@@ Introduction
DRILL (ours) 15 34 150 66
Table 4: Runtimes (in seconds) on the four datasets on GPU.
Running Time Analysis. The runtimes of DRILL and two recently proposed semi-supervised methods—including both training and inference time—on GPU are reported in Table 4. DRILL generally demonstrates higher efficiency on tabular medical data, text, and audio datasets. Although it requires more time on image datasets, this overhead is acceptable given the performance gains. Overall, DRILL offers a favorable trade-off between efficiency and effectiveness.
Conclusion and Future Work
In this paper, we propose DRILL, the very first decouple representation distillation framework for semi-supervised regression. DRIIL is implemented by the DDE, where we transform the score prediction task into the discrete buckets distribution estimation to mitigate the risk of overfitting. Then, the DDA is proposed to enhance the generalization by aligning the target buckets and non-target buckets between the teacher and student on the distribution of buckets. Extensive experimental results across multiple datasets spanning four domains—audio, image, text, and tabular medical data—demonstrate that DRILL outperforms state-of-the-art methods, indicating its superior generalization capability. Exploring more data modalities and developing a unified semi-supervised regression framework are left as directions for future work.
Appendix A Appendix
Appendix B A. Datasets
@@ Appendix A Appendix
BVCC (Cooper and Yamagishi 2021) serves as an audio quality assessment benchmark, where the objective is to predict the perceptual quality of audio samples. Annotated on a 1-to-5 scale, the labels represent averaged scores from multiple listeners. This dataset comprises 4,974 training samples, 1,066 validation samples, and 1,066 test samples. Following the protocol defined in Rankup (Huang, Fu, and Tsao 2024), we exclusively utilize the training and validation splits for performance evaluation.
Yelp (Asghar 2016) focuses on textual opinion mining, targeting the prediction of customer ratings based on reviews posted on the Yelp platform. The five-class rating labels represent distinct satisfaction levels. We employ the preprocessed Yelp data provided in the USB codebase (Huang, Fu, and Tsao 2024), consisting of 250,000 training samples, 25,000 validation samples, and 10,000 test samples. Consistent with RankUp’s methodology, only the training split is used for model training, while evaluation is performed on the validation set.
UTKFace (Zhang, Song, and Qi 2017) addresses image-based age estimation, aiming to predict the age of individuals depicted in facial images. The age labels range from 1 to 116 years. RankUp partitions this dataset into 18,964 training samples and 4,741 test samples, with experiments conducted on its aligned and cropped version.
MIMIC (Johnson et al. 2023) presents a multivariate time-series score prediction task, where the objective is to forecast SOFA scores (indicating illness severity from 0 to 24) based on patients’ clinical features over time. This dataset contains 3,623,503 training samples and 55,859 test samples.
@@
Models
$\Pi$-Model,
MeanTeacher
MixMatch UCVME CLSS RankUp DRILL
Unlabeled Batch Ratio 1.0 1.0 1.0 0.25 7.0 7.0
Regression Unlabeled Loss Ratio 0.1 0.1 0.05 - - -
Regression Unlabeled Loss Warmup 0.4 0.4 - - - -
Mixup Alpha - 0.5 - - - -
Dropout Rate - - 0.05 - - -
Ensemble Number - - 5 - - -
CLSS Lambda - - - 2.0 - -
Labeled Contrastive Loss - - - 1.0 - -
Unlabeled Contrastive Loss - - - 0.05 - -
Unlabeled Rank Loss Ratio - - - 0.01 - -
ARC Unlabeled Loss Ratio - - - - 1.0 -
ARC Loss Ratio - - - - 0.2 -
Confidence Threshold - - - - 0.95 -
Temperature - - - - 0.5 -
$\beta$ - - - - - 10
@@
Appendix C B. Competing Methods
@@
B1. Description of Baselines
$\pi$-Model (Laine and Aila 2017) achieves self-ensembling by aggregating the outputs of a single network at different training stages. Consensus predictions for unknown labels are formed through diverse regularization and augmentation strategies. Perturbations are injected into both branches, and consistency between the two perturbed outputs is enforced, ensuring a more balanced and symmetric comparison.
Mean Teacher (Tarvainen and Valpola 2017) aims to train a model via consistency between original samples and their perturbed counterparts, while using an exponential-moving-average (EMA) version of the model—termed the teacher—to supply more reliable guidance to the current student during training.
MixMatch (Berthelot et al. 2019) consolidates the dominant approaches in semi-supervised learning by generating low-entropy label guesses for augmented unlabeled examples and subsequently blending labeled and unlabeled data via MixUp.
CLSS (Dai et al. 2023) extends contrastive regression to the semi-supervised domain. It exploits a spectral ordering algorithm to extract ordinal rankings from the feature-similarity matrix of unlabeled data and then leverages these rankings as supervisory signals. The additional ranking loss enhances robustness, marking the first exploration of contrastive learning for regression tasks.
UCVME (Dai, Li, and Cheng 2023) improves training by producing high-quality uncertainty estimates for pseudo-labels via heteroscedastic regression. By explicitly modeling label uncertainty, the approach assigns greater importance to more reliable pseudo-labels. Furthermore, a novel variational model-ensembling scheme is introduced to reduce predictive noise and generate stronger pseudo-labels.
RankUp (Huang, Fu, and Tsao 2024) reframes the original regression task as a ranking problem and trains it jointly with the regression objective. By introducing an auxiliary ranking classifier whose outputs can be handled by any off-the-shelf semi-supervised classification method, RankUp seamlessly integrates existing classification techniques into regression-oriented semi-supervised learning.
@@
Strong: RandAugment (Cubuk et al. 2020)
@@
Strong: Back-Translation (Xie et al. 2020a)
@@
Appendix D C. Parameters and Data Augementation
C1. Parameter Configurations of Base Models
@@
C2. Weak and Strong Augmentation
For audio, image, and text data, we adopt the universal semi-supervised learning benchmark framework, USB (Wang et al. 2022), to implement the weak and strong augmentations following the previous work (Huang, Fu, and Tsao 2024). For the MIMIC dataset, given its time-series nature, weak augmentation involves applying random masking operations, while strong augmentation employs $m$ rounds of random masking combined with Gaussian noise injection for multivariate time-series data. Specifically, each random masking operation obscures 5% of the observed data. Each Gaussian noise application introduces additive noise sampled from a normal distribution with zero mean and a variance of 0.02. To ensure fair comparison across methods, the same augmentation strategy is applied to both DRILL and other comparative methods.
@@
Appendix E D. Additional Experimental Results
In this section, we further evaluate the effectiveness of the hyperparameters, specifically the bucket size of $L$ and $\beta$, in $\mathcal{L}_{DDA}$ using additional evaluation metrics. The results are shown in Fig.5. For the effectiveness of the bucket size $L$, compared to the $R^{2}$ metric, the MAE for DRILL exhibits greater stability under variations of $L$, see Fig.5 (a). We found that the best performance is achieved with 200 buckets for the BVCC, Yelp, and UTKFace datasets, while a lower MAE is obtained for the MIMIC dataset using 100 buckets. Regarding the effectiveness of $\beta$ in $\mathcal{L}_{DDA}$, see Fig.5 (b), the MAE values demonstrate a trend comparable to that of $R^{2}$ across all datasets, with optimal performance universally observed at $\beta=10$.
(a) DRILL model prediction example for a ground truth lable=2.
(b) DRILL model prediction example for a ground truth lable=4.
Figure 6: The visualization results of DRILL.
Appendix F E. Visualization Results
Fig. 6 presents a visualization of the prediction cases using the DRILL model, illustrating the predicted discrete buckets and their associated probability values on the Yelp dataset. As shown in Fig. 6 (a), for a regression value of 2, the Yelp dataset score prediction is made using a distributed set of LL=200 buckets that span the 0-to-5 rating range. Notably, in addition to the score of 2 (Fair/Dissatisfied) itself, the model primarily leverages the probabilities of the nearby related values of 1 (Poor/Very Dissatisfied) and 3 (Average/Neutral) for a combined prediction. A similar phenomenon can also be observed in the other example in Fig 6 (b). The results above demonstrate that DIRLL more effectively utilizes label distribution information, thereby improving SSR performance.

### v2

@@ 4.5 Efficiency Analysis
GCLSS (2025) 8 23 12 36
DKD (ours) 15 34 150 66
Table 4: Runtimes (in seconds) on the four datasets.
Running Time Analysis. The runtimes of DKD and two recently proposed semi-supervised methods, including both training and inference time on NVIDIA GeForce GTX 4060, are reported in Table 4. DKD demonstrates higher efficiency on tabular medical data, text, and audio datasets. Although it requires more time on image datasets, this overhead is acceptable given the performance gains. Overall, DKD offers a favorable trade-off between efficiency and effectiveness.
5 Conclusion
In this paper, we propose DKD, the very first dual-stream knowledge distillation framework for fully exploiting the abundance of unlabeled samples and noise mitigation in pseudo-label supervision in SSR. DKD is designed to distill both continuous-valued knowledge and distribution information, which better preserves regression magnitudes and enhances the sample efficiency. Then, the DDA is proposed to further enhance the student’s capacity to mitigate noise in pseudo-label supervision and learn a more well-calibrated regression predictor. Extensive experimental results across multiple datasets spanning four domains demonstrate that DKD outperforms state-of-the-art methods, indicating its superior generalization capability.
Appendix A Datasets
@@
BVCC Cooper and Yamagishi [2021] serves as an audio quality assessment benchmark, where the objective is to predict the perceptual quality of audio samples. Annotated on a 1-to-5 scale, the labels represent averaged scores from multiple listeners. This dataset comprises 4,974 training samples, 1,066 validation samples, and 1,066 test samples. Following the protocol defined in Rankup Huang et al. [2024], we exclusively utilize the training and validation splits for performance evaluation.
Yelp Asghar [2016] focuses on textual opinion mining, targeting the prediction of customer ratings based on reviews posted on the Yelp platform. The five-class rating labels represent distinct satisfaction levels. We employ the preprocessed Yelp data provided in the USB codebase Huang et al. [2024], consisting of 250,000 training samples, 25,000 validation samples, and 10,000 test samples. Consistent with RankUp’s methodology, only the training split is used for model training, while evaluation is performed on the validation set.
UTKFace Zhang et al. [2017b] addresses image-based age estimation, aiming to predict the age of individuals depicted in facial images. The age labels range from 1 to 116 years. RankUp partitions this dataset into 18,964 training samples and 4,741 test samples, with experiments conducted on its aligned and cropped version.
MIMIC Johnson et al. [2023] presents a multivariate time-series score prediction task, where the objective is to forecast SOFA scores (indicating illness severity from 0 to 24) based on patients’ clinical features over time. This dataset contains 3,623,503 training samples and 55,859 test samples.
@@
Models $\pi$-Model MeanTeacher MixMatch UCVME CLSS GCLSS RankUp DKD
Unlabeled Batch Ratio 1.0 1.0 1.0 1.0 0.25 0.25 7.0 7.0
Regression Unlabeled Loss Ratio 0.1 0.1 0.1 0.05 - - - -
Regression Unlabeled Loss Warmup 0.4 0.4 0.4 - - - - -
Mixup Alpha - - 0.5 - - - - -
Dropout Rate - - - 0.05 - - - -
Ensemble Number - - - 5 - - - -
CLSS Lambda - - - - 2.0 2.0 - -
Labeled Contrastive Loss - - - - 1.0 1.0 - -
Unlabeled Contrastive Loss - - - - 0.05 0.05 - -
Unlabeled Rank Loss Ratio - - - - 0.01 0.01 - -
ARC Unlabeled Loss Ratio - - - - - - 1.0 -
ARC Loss Ratio - - - - - - 0.2 -
Confidence Threshold - - - - - - 0.95 -
Temperature - - - - - - 0.5 -
The Number of Selected Features - - - - - 7 - -
$\beta$ - - - - - - - 10
@@
Appendix B Competing Methods
@@
B.1 Description of Baselines
$\pi$-Model Laine and Aila [2017] achieves self-ensembling by aggregating the outputs of a single network at different training stages. Consensus predictions for unknown labels are formed through diverse regularization and augmentation strategies. Perturbations are injected into both branches, and consistency between the two perturbed outputs is enforced, ensuring a more balanced and symmetric comparison.
Mean Teacher Tarvainen and Valpola [2017] aims to train a model via consistency between original samples and their perturbed counterparts, while using an exponential-moving-average (EMA) version of the model—termed the teacher—to supply more reliable guidance to the current student during training.
MixMatch Berthelot et al. [2019] consolidates the dominant approaches in semi-supervised learning by generating low-entropy label guesses for augmented unlabeled examples and subsequently blending labeled and unlabeled data via MixUp.
UCVME Dai et al. [2023b] improves training by producing high-quality uncertainty estimates for pseudo-labels via heteroscedastic regression. By explicitly modeling label uncertainty, the approach assigns greater importance to more reliable pseudo-labels. Furthermore, a novel variational model-ensembling scheme is introduced to reduce predictive noise and generate stronger pseudo-labels.
CLSS Dai et al. [2023a] extends contrastive regression to the semi-supervised domain. It exploits a spectral ordering algorithm to extract ordinal rankings from the feature-similarity matrix of unlabeled data and then leverages these rankings as supervisory signals. The additional ranking loss enhances robustness, marking the first exploration of contrastive learning for regression tasks.
RankUp Huang et al. [2024] reframes the original regression task as a ranking problem and trains it jointly with the regression objective. By introducing an auxiliary ranking classifier whose outputs can be handled by any off-the-shelf semi-supervised classification method, RankUp seamlessly integrates existing classification techniques into regression-oriented semi-supervised learning.
GCLSS Wang et al. [2025] extended CLSS to allow both labeled and unlabeled data to be used in the semi-supervised setting. Particularly, GCLSS uses labeled samples to regularize and rectify the recovered ordinal rankings of unlabeled data. The recovered rankings serve as pseudo-labels to supervise both the feature representations and the final predictions.
@@ B.1 Description of Baselines
Strong: RandAugment Cubuk et al. [2020]
@@
Strong: Back-Translation Xie et al. [2020]
@@ B.2 Evaluation Metric
Appendix C Parameters and Data Augementation
C.1 Parameter Configurations of Base Models
@@
C.2 Weak and Strong Augmentation
For audio, image, and text data, we adopt the universal semi-supervised learning benchmark framework, USB Wang et al. [2022], to implement the weak and strong augmentations following the previous work Huang et al. [2024]. For the MIMIC dataset, given its time-series nature, weak augmentation involves applying random masking operations, while strong augmentation employs $m$ rounds of random masking combined with Gaussian noise injection for multivariate time-series data. Specifically, each random masking operation obscures 5% of the observed data. Each Gaussian noise application introduces additive noise sampled from a normal distribution with zero mean and a variance of 0.02. To ensure fair comparison across methods, the same augmentation strategy is applied to both DKD and other comparative methods.
@@ C.2 Weak and Strong Augmentation
Appendix D Additional Experimental Results
In this section, we further evaluate the effectiveness of the hyperparameters, specifically the class size of $L$ and $\beta$, in $\mathcal{L}_{DDA}$ using additional evaluation metrics. The results are shown in Fig.5. For the effectiveness of the class size $L$, compared to the $R^{2}$, the MAE for DKD exhibits greater stability under variations of $L$, see Fig.5 (a). We found that the best performance is achieved with 200 classes for the BVCC, Yelp, and UTKFace datasets, while a lower MAE is obtained for the MIMIC dataset using 100 classes. Regarding the effectiveness of $\beta$ in $\mathcal{L}_{DDA}$, see Fig.5 (b), the MAE demonstrates a trend comparable to that of $R^{2}$ across all datasets, with optimal performance universally observed at $\beta=10$.
(a) A prediction example of DKD with ground truth $y=4$.
(b) A prediction example of DKD with $y=2$.
Figure 6: The visualization results of DKD.
Appendix E Visualization Results
Fig. 6 presents a visualization of the prediction cases using the DKD model, illustrating the predicted discrete classes and their associated probability values on the Yelp dataset. As shown in Fig. 6 (a), for a regression value of 2, the Yelp dataset score prediction is made using a distributed set of LL=200 classes that span the 0-to-5 rating range. Notably, in addition to the score of 2 (Fair/Dissatisfied) itself, the model primarily leverages the probabilities of the nearby related values of 1 (Poor/Very Dissatisfied) and 3 (Average/Neutral) for a combined prediction. A similar phenomenon can also be observed in the other example in Fig 6 (b). These results highlight DKD’s superior ability to exploit label information, which significantly enhances generalization and prevents overfitting.